# Emergent Semantics Beyond Token Embeddings: Transformer LMs with Frozen Visual Unicode Representations

### v1

@@ Abstract
INDEX TERMS emergent semantics, frozen embeddings, interpretability, language models, multilingual NLP, transformers, visual representations, Unicode
@@ 1   INTRODUCTION
The robustness of modern Large Language Models (LLMs) to orthographic variations, such as in ’I can wRiTe’, highlights a fundamental question. The model’s understanding is not derived from a single, semantically rich token for "write". Instead, it often relies on composing meaning from a sequence of semantically poor characteror subword-level tokens (e.g., ’w’, ’R’, ’i’, ’T’, ’e’). If the input vectors for these atomic units lack inherent semantic content, where does the model’s understanding originate? This paper argues and empirically demonstrates that semantics are not a property of input embeddings but an emergent phenomenon of the Transformer architecture itself.
@@
Critically, we observe that despite this radical departure from conventional practice, these “frozen-embedding” models not only converge but display robust convergence and, surprisingly, outperform architecturally identical counterparts with trainable embeddings on reasoning benchmarks like MMLU, suggesting representational interference in conventional approaches. Emergent abstraction appears within the transformer blocks themselves, not in the input vectors. This has profound implications: it implies that semantic structure is an architectural phenomenon, not an initialization artifact. Our main contributions are:
@@
The rest of this paper is organized as follows: Section II reviews related work; Section III details our method for constructing visual embeddings and Unicode tokenization; Section IV describes the experimental setup; Section V presents results and analysis; Section VI discusses implications for future architectures; and Section VII concludes. Code and all artifacts are released to foster further progress in both science and industry [^1] Repository: https://github.com/AVBochkov/Embeddings (https://huggingface.co/Bochkov).
@@ 2   RELATED WORK
Semantic representation in neural language models has received extensive attention. Early word2vec [1] and GloVe [2] models established the conceptual centrality of “meaningful” dense vector spaces, with well-documented algebraic properties in vector arithmetic. Subsequent transformer-based LMs [3] extended this paradigm to context-dependent embeddings, further entrenching the notion that trainable input vectors are the source of semantic capability. However, several lines of research challenge this view.
Character - and byte-level models, such as ByT5 [4] and CANINE [5], forego word-level tokenization for universal coverage; yet, even here, embedding matrices are trainable and morph into semantic vector spaces during pretraining. Cross-lingual and modular approaches pursue shared embedding spaces for transfer [6, 7], but rely on substantial alignment learning. Mixture-of-experts [8, 9] and adapter fusion [10] aim to modularize reasoning, yet remain dependent on trainable embeddings. Our input embedding matrix is never trained and strictly visual/surface oriented.
More radically, recent work in visual-language models [11] investigates multimodal representations by grounding language in visual contexts. Models like CLIP [12] or Flamingo [13] learn joint embeddings for text and images. However, these approaches typically combine separate, powerful encoders for each modality and aim to align two rich semantic streams, rather than enforcing a non-semantic, frozen vector space for language as we propose.
Our work differs in two key aspects: We forego data-driven embedding learning altogether, initializing the input layer with fixed, non-semantic, visual features derived from Unicode glyphs. Our primary contribution is demonstrating that high-level abstraction emerges even when the input layer is restricted to fixed, non-semantic, visual features, isolating the compositional layers as the locus of meaning.
@@ 3   METHOD
3.1.   FROZEN VISUAL UNICODE EMBEDDINGS
@@
3.2.   UNICODE-CENTRIC TOKENIZER CONSTRUCTION
@@
The core approach is based on bijective mapping of Unicode codepoints to individual tokens, whenever possible, ensuring a 1:1 relationship between characters and tokens across a wide range of scripts. Reserved, surrogate, and private Unicode ranges are systematically exploited to store custom and multi-character tokens, allowing for efficient representation of commonly occurring substrings and entire words—particularly beneficial for high-frequency terms shared across models and languages.
The tokenizer’s hybrid design draws from frequent multi-character patterns identified in top-performing, state-of-the-art models: approximately 19,000 such tokens are incorporated. This structure enables robust handling of diverse languages and scripts, mitigating the fragmentation common in byte-pair encoding (BPE) and unigram tokenizers—especially for underrepresented scripts and code-mixed text.
@@
Furthermore, since all token indices and their corresponding Unicode/codepoint representations are deterministic and exhaustively specified, this methodology enables the direct creation of a pre-initialized embedding matrix of shape vocab_size × n_embed. Such a matrix can be either used as a fixed (non-trainable) embedding or as a high-quality cold start, reducing the need for extensive embedding pre-training. Combining strict Unicode mapping with multi-character tokens, the bvv241 family achieves highly stable performance for Asian, Arabic, and underrepresented scripts—where differences from SOTA are minimal. For Indo-European (Latin and partially Cyrillic) languages, compression is less aggressive, but the model gains in predictability, recoverability, and linguistic inclusivity.
$3.$   MODEL ARCHITECTURE
We employ a standard decoder-only Transformer architecture, similar to GPT-2. All models use a hidden dimension d_model=1024, multi-head attention, and GELU activations. The key modification is that the token embedding layer (nn.Embedding) has its weights frozen to the precomputed visual matrix. For comparison, we train baseline models with identical architectures but with a standard, randomly initialized and trainable embedding layer.
@@ 4   EXPERIMENTAL SETUP
4.1.   TRAINING DATASETS
The training corpus comprises a multilingual mix including subsets of Wikipedia (en, ru, zh), and SFT datasets (10% insertion in pretrain) SQuAD-v2, TriviaQA, HotpotQA, NQ-Open, BoolQ, CommonsenseQA, and ARC, totalling approximately 9B tokens. Wikipedia data constitutes roughly 5% of its total size.
Model Name Parameters and Vocab Size Training Corpus and Embedding Type
pro_bvv_en 0.2 B, 65536 English, Frozen
pro_bvv_unfrozen 0.2 B, 65536 English, Trainable
pro_bvv_ru 0.2 B, 65536 EN+Russian, Frozen
pro_bvv_zh 0.2 B, 65536 EN+Chinese, Frozen
best_bvv_ru 0.5 B, 131072 EN+Russian, Frozen
best_bvv_unfrozen_ru 0.5 B, 131072 EN+Russian, Trainable
best_bvv_zh 0.5 B, 131072 EN+Chinese, Frozen
best_bvv_unfrozen_zh 0.5 B, 131072 EN+Chinese, Trainable
max_bvv_ru 0.4 B, 131072 EN+Russian, Frozen
max_bvv_zh 0.4 B, 131072 EN+Chinese, Frozen
nemo_bvv_ru 0.4 B, 131072 EN+Russian, Frozen
nemo_bvv_zh 0.4 B, 131072 EN+Chinese, Frozen
Table 1: Model Parameters and Vocabulary Statistics. Overview of models used in experiments. ’Frozen’ models use our fixed visual embeddings. ’Trainable’ models are baselines with standard, randomly initialized, trainable embeddings. All other architectural parameters (d_model=1024, layers, etc.) are held constant within each size class (0.2B, 0.4B, 0.5B) between frozen and trainable embeddings mode. The prefixes ’pro’, ’max’, and ’best’ correspond to models with 0.2B, 0.4B, and 0.5B parameters, respectively.
4.2.   BASELINES AND COMPARISONS
Baselines include classic transformer LMs with trainable embedding layers matched for architecture, parameter count, and tokenizer coverage. Our primary comparisons are between models with ’Frozen’ embeddings and their ’Trainable’ counterparts (e.g., ’best_bvv_ru’ vs. ’best_bvv_unfrozen_ru’), which share identical architecture and training data.
4.3.   EVALUATION
@@
$4.$   IMPLEMENTATION DETAILS
Training was performed on H100 80 Gb with batch sizes up to 16 and block sizes of 1024. Gradient accumulation and mixed precision were used for efficiency.
@@ 5   RESULTS
Before presenting our results, it is crucial to frame their context. This study does not aim to achieve state-of-the-art performance on benchmarks like MMLU. Our models are intentionally constrained in scale (up to 0.5B parameters) and trained on a modest 9B token dataset. The primary objective is to demonstrate the feasibility of learning with non-semantic embeddings and to compare the learning dynamics between our proposed frozen-embedding models and traditional trainable-embedding baselines under identical conditions. The following metrics should therefore be interpreted as evidence of emergent abstraction, not as a challenge to large-scale models.
5.1.   MODEL CONVERGENCE
@@
A primary concern was whether a model with non-semantic, frozen embeddings could converge at all. Figure 2 shows the training loss curves for a 0.5B parameter frozen-embedding model (best_bvv_ru) and Figure 3 shows its trainable-embedding counterpart (best_bvv_unfrozen_ru). Both models exhibit stable convergence profiles, with nearly identical loss trajectories. This demonstrates that the Transformer architecture is capable of learning effectively even when the input layer provides only fixed structural information.
Figure 2: Learning curves (best_bvv_ru) for frozen embedding.
Figure 3: Learning curves (best_bvv_unfrozen_ru) for trainable embedding.
5.2.   PERFORMANCE OF FROZEN VS. TRAINABLE EMBEDDING MODELS
@@
Our primary finding challenges the conventional wisdom that trainable, semantically-rich embeddings are necessary for high-level reasoning. As shown in Figure 4 and Figure 5, while our frozen-embedding models demonstrate robust convergence across all tasks, they show a striking and consistent performance advantage on the MMLU reasoning benchmark compared to their architecturally identical counterparts with trainable embeddings.
@@
Figure 4: Performance Comparison: MMLU sub tasks with score greater than 27%.
@@
Figure 5: Performance Comparison: Frozen vs. Trainable Embedding Models.
@@
For instance, best_bvv_ru (0.5B params, frozen embeddings) achieves an MMLU score of 22.29, whereas best_bvv_unfrozen_ru (0.5B params, trainable embeddings) scores only 11.37 — a nearly 2x performance gap. This pattern holds across different model sizes and languages.
@@
We hypothesize this phenomenon, which we term "representational interference," stems from the dual burden placed on trainable embeddings. A trainable embedding layer must simultaneously learn low-level orthographic features (e.g., character shapes, token boundaries) and high-level semantic abstractions within the same parameter space. This creates a conflict during optimization, leading to a suboptimal representation that is neither a perfect structural encoder nor a pure semantic vector.
@@
In contrast, our frozen visual embeddings provide a stable, information-rich structural foundation. By offloading the task of representing token form to a fixed, deterministic layer, we free the Transformer’s attention and feed-forward layers to focus exclusively on their core competency: composing these structural primitives into high-level meaning. The model does not waste capacity learning what a token looks like; it only learns what to do with it.
@@
5.3.   EMERGENT SEMANTICS VISUALIZATION OF INPUT TOKEN EMBEDDINGS

### v4

@@ 1 INTRODUCTION
The robustness of modern Large Language Models (LLMs) to orthographic variations, such as in ’I can wRiTe’, highlights a fundamental question. The model’s understanding is not derived from a single, semantically rich token for "write". Instead, it often relies on composing meaning from a sequence of semantically poor character- or subword-level tokens (e.g., ’w’, ’R’, ’i’, ’T’, ’e’). If the input vectors for these atomic units lack inherent semantic content, where does the model’s understanding originate? This paper argues and empirically demonstrates that semantics are not a property of input embeddings but an emergent phenomenon of the Transformer architecture itself.
@@
Critically, we observe that despite this radical departure from conventional practice, these “frozen-embedding” models not only converge but display robust convergence and, surprisingly, outperform architecturally identical counterparts with trainable embeddings on reasoning benchmarks like MMLU, suggesting representational interference in conventional approaches. This suggests that emergent abstraction appears within the transformer blocks themselves, implying that semantic structure is an architectural phenomenon, not an initialization artifact. Our main contributions are:
@@
The rest of this paper is organized as follows: Section II reviews related work; Section III details our method for constructing visual embeddings and Unicode tokenization; Section IV describes the experimental setup; Section V presents results and analysis; Section VI discusses implications for future architectures; and Section VII concludes. Code and all artifacts are released to foster further progress in both science and industry. [^1] Repository: https://github.com/AVBochkov/Embeddings ( https://huggingface.co/Bochkov ).
@@ 2 RELATED WORK
Semantic representation in neural language models has received extensive attention. Early word2vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) models established the conceptual centrality of “meaningful” dense vector spaces, with well-documented algebraic properties in vector arithmetic. Subsequent transformer-based LMs (Radford et al., 2019) extended this paradigm to context-dependent embeddings, further entrenching the notion that trainable input vectors are the source of semantic capability. However, several lines of research challenge this view.
Character- and byte-level models, such as ByT5 (Xue et al., 2022) and CANINE (Clark et al., 2022), forego word-level tokenization for universal coverage; yet, even here, embedding matrices are trainable and morph into semantic vector spaces during pretraining. Cross-lingual and modular approaches pursue shared embedding spaces for transfer (Lample et al., 2018; Lample and Conneau, 2019), but rely on substantial alignment learning. Mixture-of-experts (Shazeer et al., 2017; Fedus et al., 2022) and adapter fusion (Pfeiffer et al., 2020) aim to modularize reasoning, yet remain dependent on trainable embeddings. Our input embedding matrix is never trained and strictly visual/surface oriented.
More radically, recent work in visual-language models (Tan and Bansal, 2022) investigates multimodal representations by grounding language in visual contexts. Models like CLIP (Radford et al., 2021) or Flamingo (Alayrac et al., 2022) learn joint embeddings for text and images. However, these approaches typically combine separate, powerful encoders for each modality and aim to align two rich semantic streams, rather than enforcing a non-semantic, frozen vector space for language as we propose.
The idea of incorporating visual information from text glyphs is not entirely new. Prior work has successfully utilized visual features to enhance specific tasks. For instance, visual representations have been used to improve machine translation, especially for handling rare words and achieving robust cross-lingual transfer Salesky et al. (2023); Wang et al. (2020). Other studies have leveraged glyph-aware embeddings for better representation of logographic languages like Chinese Dai and Cai (2017) or to learn character-level compositionality Liu et al. (2017).
Our work differs from these approaches in two fundamental aspects. First, whereas prior studies use visual features as a trainable component or an auxiliary signal to enrich semantic embeddings, our methodology employs them as a completely frozen, non-semantic substrate. Second, our primary goal is not to improve a specific downstream task, but to use this constraint as a tool to investigate a foundational question: whether semantic understanding can emerge exclusively within the Transformer’s compositional layers, without any semantic information provided at the input. In our work, this reframes the role of visual features from an enhancement to a basic, structural primitive.
@@ 3 METHOD
3.1 FROZEN VISUAL UNICODE EMBEDDINGS
@@
To investigate the influence of PCA and normalization, as well as to understand the fundamental significance of visual representations of token texts, a maximally simplified approach was implemented. For a vocabulary of 65,536 tokens, a frozen pre-computed tensor was constructed based on binary components of a 16-dimensional vector for each token—essentially, each token was represented by its index as a 16-dimensional vector where each component is a single bit, thereby eliminating the influence of PCA algorithms, normalization, and additional information about the visual representation of the token text (model ’Model_16_bit’). Similarly, to study the influence of PCA and normalization algorithms separately, another pre-computed tensor of 16-dimensional vector representations was constructed, but this time with the application of PCA and normalization (model ’Model_16_float’)
For ablation study purposes, three additional variants with binary components based on noise were created for 64-dimensional, 256-dimensional, and 1024-dimensional vectors. Three corresponding variants with applied PCA and normalization were also developed to investigate the impact on convergence and metrics in the absence of information about the visual representation of the token text itself. (’Model_1024_bit’, ’Model_1024_float’, ’Model_256_bit’, ’Model_256_float’, ’Model_64_bit’, ’Model_64_float’).
3.2 UNICODE-CENTRIC TOKENIZER CONSTRUCTION
@@
3.2.1 N-gram Extraction
The first part is n-grams extraction: automatic extraction of the most frequent n-grams (sequences of n characters) from Wikipedia and Wiki-source text corpora across 32 world languages. For each language, the algorithm extracts n-grams of four different sizes (n = 2, 3, 4, 5 characters), counts their frequency using the Counter data structure, and preserves the top 1000 most frequent n-grams for each value of n. Determination of whether a character belongs to CJK languages is performed by checking if it falls within Unicode ranges: U+4E00–U+9FFF for Chinese ideographs, U+3040–U+309F for Hiragana, U+30A0–U+30FF for Katakana, and U+AC00–U+D7AF for Korean Hangul. Results are saved in a structured format containing the language code, its full name, and a dictionary with top n-grams for each length, enabling subsequent use of this data for language identification tasks, morphological pattern analysis, or language model creation.
3.2.2 Tokenizer Implementation and Vocabulary
The second part is tokenizer implementation: algorithm implements the creation of a multilingual tokenizer with a fixed vocabulary size of 65,536 tokens, based on statistically significant n-grams from the previous stage. The algorithm employs a direct mapping strategy for Unicode characters in the range 0x0000–0xD7FF (55,296 positions) to corresponding token indices, ensuring complete coverage of the Unicode Basic Multilingual Plane. The most frequent n-grams (8,177 unique sequences in total) are distributed across two ranges: the first 2,048 n-grams are placed in the Unicode surrogate pairs zone (0xD800–0xDFFF), which is not used for regular characters, while the remaining ones occupy the range 0xE100–0xF8F5. Special tokens, including sequence start and end markers, dialogue markup tokens, and technical markers, are reserved in the range 0xE000–0xE0FF (256 positions). Unused positions are filled with dummy tokens to achieve the target vocabulary size of 65,536 elements.
So, the core approach is based on bijective mapping of Unicode codepoints to individual tokens, whenever possible, ensuring a 1:1 relationship between characters and tokens across a wide range of scripts. Reserved, surrogate, and private Unicode ranges are systematically exploited to store custom and multi-character tokens, allowing for efficient representation of commonly occurring substrings and entire words—particularly beneficial for high-frequency terms shared across models and languages.
3.2.3 Vocabulary Expansion
The next version or expansion of the tokenizer from 65,536 to 131,072 tokens represents a qualitative transition from basic Unicode coverage to a hybrid architecture capable of efficiently processing modern multilingual corpora and specialized domains. In the base 65,536-token version, virtually the entire vocabulary is occupied by direct Unicode character mapping (approximately 55,000 positions) and statistically identified n-grams (about 8,000), leaving minimal space for additional optimizations. Doubling the vocabulary size to 131,072 preserves all advantages of bijective Unicode mapping while simultaneously integrating approximately 19,000 high-frequency multi-character tokens extracted from analysis of leading language models. These additional tokens include whole words, stable collocations, technical terms, and programming constructs that frequently appear across various languages and domains. This approach dramatically reduces tokenization fragmentation for specialized texts, code, and low-resource languages, where traditional BPE tokenizers may split semantically coherent units into numerous small fragments.
Architecturally, the extended tokenizer utilizes the additional address space 0x10000–0x1FFFF to accommodate these high-frequency patterns, systematically exploiting reserved and private Unicode ranges. This achieves an optimal balance between universality (every Unicode character is guaranteed to have a token) and efficiency (frequent sequences are encoded with a single token). The hybrid design is particularly effective for processing mixed texts containing code-switching, technical documentation with embedded code snippets, and texts in non-Latin scripts, where standard tokenizers often demonstrate suboptimal performance due to bias toward English-language corpora during training.
@@ 3.2.3 Vocabulary Expansion
Furthermore, since all token indices and their corresponding Unicode/codepoint representations are deterministic and exhaustively specified, this methodology enables the direct creation of a pre-initialized embedding matrix of shape vocab_size × n_embed. Such a matrix can be either used as a fixed (non-trainable) embedding or as a high-quality cold start, reducing the need for extensive embedding pre-training. Combining strict Unicode mapping with multi-character tokens, the bvv241 family achieves highly stable performance for Asian, Arabic, and underrepresented scripts—where differences from SOTA are minimal. For Indo-European (Latin and partially Cyrillic) languages compression is less aggressive.
3.3 MODEL ARCHITECTURE
We employ a standard decoder-only Transformer architecture, similar to GPT-2. All models use a hidden dimension d_model=1024, n_layer = 16, multi-head attention, n_head = 32, rotary embeddings, and GELU activations. The key modification is that the token embedding layer (nn.Embedding) has its weights frozen to the precomputed visual matrix (labeled as ’Model UNI_GLIF’, total parameters: 335.8M, trainable parameters: 268.7M, frozen parameters: 67.1M). For comparison, we train baseline model with identical architectures but with a standard, randomly initialized and trainable embedding layer (labeled as ’Model unfrozen’, total parameters: 335.8M, trainable parameters: 335.8M, frozen parameters: 0.0M). For ablation study: model with 16-dimensional binary frozen embeddings (labeled as ’Model_16_bin’, total parameters: 269.7M, trainable parameters: 268.7M frozen parameters: 1.0M, d_model=1024, n_embed=16), model with 16-dimensional float frozen embeddings (labeled as ’Model_16_float’, total parameters: 269.7M, trainable parameters: 268.7M frozen parameters: 1.0M, d_model=1024, n_embed=16). The 16-dimensional vector is projected into the model’s 1024-dimensional space via repetition (repeat_interleave(64, dim=-1)), creating a simple, non-trainable mapping.). The model does not use a tied input/output embedding design; the output projection head is trained independently.
@@ 4 EXPERIMENTAL SETUP
4.1 TRAINING DATASETS
The training corpus comprises a multilingual mix including subsets of Wikipedia (en, ru, zh), and SFT datasets (10% of iterations) SQuAD-v2, TriviaQA, HotpotQA, NQ-Open, BoolQ, CommonsenseQA, and ARC: approximately 4B tokens (train 4’121’073’286 tokens + SFT train 419’136 tokens, validation 457’897’032 tokens).
4.2 BASELINES AND COMPARISONS
Baselines include classic transformer LMs with trainable embedding layers matched for architecture, parameter count, and tokenizer coverage. Our primary comparisons are between models with ’Frozen’ embeddings and their ’Trainable’ counterparts (e.g. ’Model unfrozen’ vs. ’Model UNI_GLIF’), which share identical architecture and training data and tokenizer ’bvv241-2-3’ (vocab size: 65536).
4.3 EVALUATION
@@
4.4 IMPLEMENTATION DETAILS
Training was performed on H100 80 Gb with batch sizes up to 18 and block sizes of 1024. Evaluation interval was 500 (each 500 iterations MMLU, ARC, C-SENSE metrics recalculated), gradient accumulation steps parameter was 250, total iterations was set to 400’000 due to strict computation and time limits (1.7 epochs). The final model checkpoint after 400’000 iterations was used for final evaluation.
@@ 5 RESULTS
Before presenting our results, it is crucial to frame their context. This study does not aim to achieve state-of-the-art performance on benchmarks like MMLU. Our models are intentionally constrained in scale (up to 0.3B parameters) and trained on a small dataset. The primary objective is to demonstrate the feasibility of learning with non-semantic embeddings and to compare the learning dynamics between our proposed frozen-embedding models and traditional trainable-embedding baselines under identical conditions. The following metrics should therefore be interpreted as evidence of emergent abstraction, not as a challenge to large-scale models.
5.1 MODEL CONVERGENCE
@@
A primary concern was whether a model with non-semantic, frozen embeddings could converge at all. Figure 2 and Figure 3 show the training loss curves and MMLU, ARC-e performance for a 0.3B frozen-embedding model(’Model UNI_GLIF’) and its trainable-embedding counterpart (’Model unfrozen’). Both models exhibit stable convergence profiles, with nearly identical loss trajectories.
Of particular note is the convergence of models with 16-dimensional binary embeddings (Model_16_bit) and 16-dimensional float embeddings (Model_16_float), which effectively carried information only about the token index to distinguish one token from another, without any information about the visual representation of the token text.
Figure 2: Learning curves for frozen vs. trainable embedding models.
Figure 3: MMLU and ARC-e curves for frozen vs. trainable embedding models.
5.2 PERFORMANCE OF FROZEN VS. TRAINABLE EMBEDDING AND SOTA MODELS
@@
Our primary finding challenges the conventional wisdom that trainable, semantically-rich embeddings are necessary for high-level reasoning. As shown in Figure 4 and Figure 5, while our frozen-embedding models demonstrate robust convergence across all tasks, they show a striking and consistent performance advantage on the MMLU reasoning benchmark compared to their architecturally identical counterparts with trainable embeddings. We also conducted a comparison with state-of-the-art (SOTA) models SmolLM (Allal et al., 2024) and SmolLM2 (Allal et al., 2025) (360M and 135M) that have a comparable number of parameters.
@@
Figure 4: Performance Comparison: MMLU sub tasks with score greater than 25%.
@@
Figure 5: Performance Comparison: Frozen vs. Trainable Embedding and SOTA SmolLM Models.
@@
For instance, ’Model UNI_GLIF’ (0.3B params, frozen embeddings) achieves an MMLU score of 23.81, whereas ’Model unfrozen’ (0.3B params, trainable embeddings) scores only 12.54 — a nearly 2x performance gap.
@@
We hypothesize this phenomenon, which we term "representational interference," stems from the dual burden placed on trainable embeddings. A trainable embedding layer must simultaneously learn low-level token identification task in terms of vector representation and high-level semantic abstractions within the same parameter space. This creates a conflict during optimization, leading to a suboptimal representation that is neither a perfect structural encoder nor a pure semantic vector.
@@
In contrast, our frozen visual embeddings provide a stable, information-rich structural foundation. By offloading the task of representing token form to a fixed, deterministic layer, we free the Transformer’s attention and feed-forward layers to focus exclusively on their core competency: composing these structural vector primitives into high-level meaning. The model does not waste capacity learning what a token looks like; it only learns what to do with it.
@@
It is critical to contextualize our baseline’s performance against the SOTA models used in our comparisons. The SmolLM models (>28% MMLU) were trained on up to 600B tokens (smollm-corpus), whereas our experiments were deliberately constrained to a 4B token dataset—a  150x difference in compute and data. Our goal was not to match SOTA but to ensure a perfectly fair comparison between our proposed method and the baseline under identical, resource-constrained conditions (the only difference was the frozen/unfrozen embedding layer).
Under these severe data constraints, it is the conventional trainable baseline (Model_unfrozen, 12.5% MMLU) that dramatically underperforms. In contrast, our frozen-embedding model (Model_UNI_GLIF, 23.8% MMLU) achieves a score comparable to established models like GPT-2 137M (25.8%-26.29% MMLU).
5.3 EMERGENT SEMANTICS VISUALIZATION OF INPUT TOKEN EMBEDDINGS AND DEEP TRANSFORMER LAYERS